OpenStudy (anonymous):

What is the equation of the line passing through (0, 2) and perpendicular to the line y = 4x – 3 in slope-intercept form?

OpenStudy (lgbasallote):

do you know the slope of your equation?

OpenStudy (anonymous):

The slope of a perpendicular line is the opposite reciprocal of 4

OpenStudy (anonymous):

one of these ore the answer y = x + 2 y = x + 2 y = 4x + 2 y = –4x + 3

OpenStudy (anonymous):

Ok: So the slope will be -1/4 y = -1/4x + b 2 = -1/4(0) + b 2 = 0 + b b = 2 y = -1/4x + 2